Circular No. 8389

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/2004 S 1 AND S/2004 S 2 C. C. Porco, CICLOPS, Space Science Institute, Boulder; and the

Cassini Imaging Science Team (cf. Porco et al. 2004, Science, in preparation) report the discovery of two new satellites of Saturn, S/2004 S 1 and S/2004 S 2, both orbiting between Saturn I (Mimas) and Saturn II (Enceladus). Revolution periods, orbital semimajor axes, and estimated satellite diameters are as follows: S/2004 S 1, 1.01 days, 194000 km, 3 km; S/2004 S 2, 1.14 days, 211000 km, 4 km. There is a slight possibility that S/2004 S 1 is identical with S/1981 S 14, observed in a single Voyager image on 1981 Aug. 23 (IAUC 6162), but this is still under investigation. Both orbits appear to be approximately circular and equatorial. Estimated diameters assume disk-integrated reflectivities of 0.15 at a phase angle of 67^{o}. Visual magnitudes seen from the earth are estimated to be in the range of ∼ 22– 23. The objects were first discovered in imaging sequences taken on 2004 June 1 that were specifically designed to search for new satellites in the inner saturnian system. In these sequences, S/2004 S 1 was observed over 6.0 hr, while S/2004 S 2 was observed over 9.3 hr. All exposures were 4.6 s and taken with the narrow-angle camera through the clear filters (Porco et al. 2004, Space Sci. Rev., in press). S/2004 S 1 was subsequently found in other Cassini narrow-angle camera images taken three weeks earlier with exposure times of 18 s.

(854) FROSTIA R. Behrend, Geneva Observatory; L. Bernasconi, Les Engarouines,

France; A. Klotz, TAROT, Observatoire de Haute-Provence; and R. Durkee, Minneapolis, MN, report that photometric observations of (854), obtained during July 16–19, show a regularly shaped light curve with period 1.565 day and amplitude 0.33 mag, with sharp attenuations characteristic of mutual-eclipse/occultation events (observed at minima of the regular light curve) having depth ∼ 0.7 mag and duration ∼ 3.7 hr — strongly suggesting that (854) is a binary system with properties similar to those of (1089), (1313), and (4492). The first complete event occurred on July 16.92 UT.

SUPERNOVA 2004dj IN NGC 2403 Visual magnitude estimates: Aug. 7.899 UT, 11.8 (R. J. Bouma, Gron-

ingen, The Netherlands); 10.89, 12.1 (A. Mizser, Budapest, Hungary); 13.916, 11.9 (K. Hornoch, Lelekovice, Czech Republic).

2004 August 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green